OpenStudy (anonymous):

Find the solution of 5x-y=11 3x-2y=8

OpenStudy (zzr0ck3r):

5x-y=11 implies y=5x-11 plug this into the other equation 3x-2y=8 for y so 3x-2(5x-11)=8

OpenStudy (zzr0ck3r):

can you solve for x?

OpenStudy (anonymous):

x=2